# Towards Quantum Universal Hypothesis Testing thanks: B. Chen would like to acknowledge the support of a Google gift fund

### v1

@@ Abstract
Index Terms:
Quantum Hypothesis Testing, Universal Hypothesis Testing, Quantum Universal Hypothesis Testing, Quantum State Tomography.
@@ I Introduction
This parallel - both in problem formulations and in solutions - between classical and QHT is quite apparent in many of the results in the literature. Besides the Quantum Stein’s lemma[5], we can use the Chernoff bound [7] when both types of errors are considered in a framework that differs from the Neyman-Pearson setup [8]. Indeed, this observation is what motivates the present work: although there is a line of work focused on it in classical hypothesis testing, the so-called universal hypothesis testing (UHT) has remained largely unexplored in the quantum setting. UHT addresses the problem of determining whether a sequence of samples is generated from a known nominal distribution. It is, in essence, equivalent to the goodness of fit testing [9, 10, 11, 12]. Hoeffding’s formulation for UHT [13], in which the unknown alternative is a fixed distribution in the probability simplex, leads to a surprising result: even though this alternative is completely unknown, the achievable performance (in terms of the type II error probability, or more precisely, its exponential decay rate) is essentially the same as that of simple hypothesis testing when the alternative is fully specified.
@@
While QUHT has not been formally defined or studied to date, there have been several related works reported in the quantum literature. These include extensions of Sanov’s theorem [14, 15, 16, 17], composite hypothesis testing [18], and Hoeffding-type error exponents in binary quantum testing [19, 20, 15]. Additionally, there are several lines of research in quantum circuits and information where QUHT can play an important role. Examples include quantum circuit equivalence [21, 22], quantum state discrimination [23, 24, 25, 26], and quantum state verification and certification [27, 28, 29].
The UHT formulation is often referred to as the one-sample test, especially in the machine learning (ML) community. One can similarly define the quantum analog to the two-sample test [30] in which two sequences of samples (quantum systems) are given and the task is to determine whether the two sequences are generated from the same distribution (quantum state). Our proposed solution to the UHT can be easily generalized to the two-sample setting, since respective concentrations of the two-sample sequences will ensure exponential consistency of the hypothesis test when QST is applied to both sequences.
Figure 1: QST ensures the concentration of the reconstructed state around the true state $\sigma$. When $\sigma\neq\rho$, separation of the reconstructed state from the nominal state $\rho$ is ensured asymptotically.
@@ II Problem Setup
Let $\mathcal{H}$ be a finite-dimensional Hilbert space. We denote by $\mathcal{L}(\mathcal{H})$ the space of linear operators on $\mathcal{H}$ and by $\mathcal{U}(\mathcal{H})\subset\mathcal{L}(\mathcal{H})$ the group of unitary operators satisfying $U^{\dagger}U=UU^{\dagger}=I_{\mathcal{H}}$. The set of density operators $\mathcal{D}(\mathcal{H})\subset\mathcal{L}(\mathcal{H})$ consists of positive semidefinite, unit-trace operators [31]. In the special case $\dim(\mathcal{H}_{2})=2$, $\mathcal{D}(\mathcal{H}_{2})$ is the set of qubit density operators. We write $\operatorname{Tr}(\cdot)$ for the trace functional, and for any $\sigma\in\mathcal{D}(\mathcal{H})$, its $m$-fold tensor product is denoted $\sigma^{\otimes m}\in\mathcal{D}(\mathcal{H}^{\otimes m})$. Throughout, $\log$ and $\ln$ denote logarithms base 2 and $e$, respectively. The trace norm of $\rho\in\mathcal{L}(\mathcal{H})$ is $\|\rho\|_{1}=\operatorname{Tr}[\sqrt{\rho^{\dagger}\rho}]$, and the trace distance between $\rho$ and $\sigma$ is $T(\rho,\sigma)=\tfrac{1}{2}\|\rho-\sigma\|_{1}$. The quantum fidelity is $F(\rho,\sigma)=\|\sqrt{\rho}\,\sqrt{\sigma}\|_{1}^{2}$, and the quantum relative entropy is $D(\rho\|\sigma)=\operatorname{Tr}[\rho\ln\rho-\rho\ln\sigma]$ [31].
@@ II-A Quantum Hypothesis Testing
$H_{0}:\sigma$ $=\rho_{0}$ (1)
$H_{1}:\sigma$ $=\rho_{1}.$ (2)
Any test is specified by a decision rule $M_{m}:\mathcal{D}(\mathcal{H})^{m}\to\{0,1\}$, such that $\hat{H}=H_{0}$ if $M_{m}(\sigma^{\otimes m})=0$, and $\hat{H}=H_{1}$ if $M_{m}(\sigma^{\otimes m})=1$. Here we use $\hat{H}$ to denote the declared hypothesis and will reserve $H$ for the true hypothesis. With this notation, we can now define the type I and type II error probabilities as follows:
$\alpha(M_{m})$ $:=\mathbb{P}_{M_{m}}[\hat{H}=H_{1}|H=H_{0}]$ (3)
$\beta(M_{m})$ $:=\mathbb{P}_{M_{m}}[\hat{H}=H_{0}|H=H_{1}].$ (4)
When it is clear from context, we may denote them by $\mathbb{P}[H_{1}|H_{0}]$ and $\mathbb{P}[H_{0}|H_{1}]$, respectively. Here use the term symmetric to describe the setting where we attempt to minimize the sum of errors $\alpha(M_{m})+\beta(M_{m})$, and asymmetric to denote the setting where we restrict the type I error to be less than $\alpha$ and we minimize the type II error.
With a single (or finite) copy of the quantum systems, the Helstrom bound [32] [33] provides a lower bound on the sum of the two error probabilities (which is in essence the average error probability, if an equal prior is assumed):
$\alpha(M_{m})+\beta(M_{m})\geq\frac{1}{2}\left(1-\frac{1}{2}\left\|\rho^{\otimes m}-\sigma^{\otimes m}\right\|_{1}\right).$ (5)
The Quantum Stein’s lemma [4, 5], on the other hand, provides an asymptotic performance bound when a large number of copies are available. Specifically, subject to a constraint on the type I error probability, there exists a sequence of decision rules ${M_{m}}$ such that:
$\lim_{m\to\infty}\frac{1}{m}\ln(\beta_{\rho_{0},\rho_{1}}(M_{m}))=-D(\rho_{0}\|\rho_{1}),$ (6)
Additionally, no other test achieves a better exponential decay rate.
@@ II-B Quantum Universal Hypothesis Testing
QUHT is a natural extension of the classical universal hypothesis testing framework to the quantum setting. Here, the null hypothesis asserts that the quantum systems are specified by a known nominal state $\rho$, while the alternative states otherwise, i.e.,
$H_{0}:\sigma$ $=\rho$ (7)
$H_{1}:\sigma$ $\neq\rho.$ (8)
This formulation directly mirrors Hoeffding’s formulation of the classical UHT, where the alternative is represented by an unknown but fixed state $\sigma$ that is different from $\rho$. For a given decision rule $M_{m}$, we denote the type I and II error probabilities as $\alpha(M_{m})$ and $\beta_{\rho,\sigma}(M_{m})$, respectively. We note that while any decision rule should not depend on $\sigma$ as it is unknown, the actual performance of the rule (in particular, type II error probability) is typically a function of $\sigma$. We may occasionally use $\alpha$ to refer to a constraint on the type I error, as well as the function describing the type I error of a decision rule, with the distinction to be clarified by context.
The generality of the QUHT formulation lies in the generality of the alternative hypothesis. While this motivates its application to broad applications in quantum systems, it also presents unique challenges. Many techniques developed for QHT rely on knowledge of both states — e.g., optimal tests for symmetric or structured alternatives [34, 14, 16] - and are thus not directly applicable when the alternative is entirely unspecified. An exception is found in recent work [17], which constructs a "quantum empirical distribution" and proves large-deviation results reminiscent of those in the classical setting. We further discuss its connection to our work in Section IV.
@@ II-C Quantum Two-Sample Test
$H_{0}:\sigma$ $=\rho$ (9)
$H_{1}:\sigma$ $\neq\rho.$ (10)
While the form appears identical to the one-sample test (i.e., QUHT), the setup differs in one key way. The two-sample test has no knowledge of both $\rho$ and $\sigma$ and is given two sequences of samples, whereas the one-sample test has perfect knowledge of $\rho$ and is given a single sequence generated from the unknown state $\sigma$.
@@ III Results
This section presents our main results on QUHT for both the one- and two-sample settings. As alluded to in Section I, we primarily leverage the concentration results found in the QST literature to obtain the error exponent for type II error probabilities. The only exception is when the nominal state is pure, in which case we obtain an optimal error exponent.
Table I presents our main results for both one-sample and two-sample variants of QUHT. To emphasize the dominant error-exponent terms, we have omitted those that vanish asymptotically as the number of copies of the quantum states under test grows without bound. Here, $m$ and $n$ denote the numbers of identical copies of the states $\rho$ and $\sigma$, respectively, as introduced in Section II-C. Throughout our analysis, the type I error threshold $\alpha\in(0,1)$ is treated as a fixed constant. In the two-sample setting, we restrict attention to the regime in which $m$ and $n$ increase at the same rate and define $k=\min\{m,n\}$ to express all bounds in terms of the effective sample size.
@@
Theorem 1
Indep. $[F(\rho,\sigma)]^{m}$
@@
Theorem 2
Indep. Lemma 1 $\exp\left\{-\frac{m\left\|\rho-\sigma\right\|_{1}^{2}}{54}\right\}$
@@
Theorem 3
Indep. Lemma 1 $\exp\left\{-\frac{m\left\|\rho-\sigma\right\|_{1}^{2}}{216}\right\}$
@@
Theorem 4
Indep. Lemma 2 $\exp\left\{-\frac{m\left\|\rho-\sigma\right\|_{1}^{2}}{86\cdot d^{3}}\right\}$
@@
Theorem 5
Indep. Lemma 2 $\exp\left\{-\frac{k\left\|\rho-\sigma\right\|_{1}^{2}}{344\cdot d^{3}}\right\}$
@@
Theorem 6
Entangled Lemma 3 $\exp\left\{-\frac{m\left\|\rho-\sigma\right\|_{1}^{2}}{2}\right\}$
@@
Theorem 7
Entangled Lemma 3 $\exp\left\{-\frac{k\left\|\rho-\sigma\right\|_{1}^{2}}{8}\right\}$
@@ III-A Pure State QUHT
Take $\rho=\lvert\phi\rangle\langle\phi\rvert\in\mathcal{D}(\mathcal{H})$ a pure state density operator and $\sigma\in\mathcal{D}(\mathcal{H})$ a (potentially mixed) density operator. In the QUHT setting, given $\sigma^{\otimes m}$ and $\alpha\in(0,1)$, using independent measurements there exists a decision rule $M_{m}$, such that
@@
$\beta_{\rho,\sigma}(M_{m})$ $=\left[F(\rho,\sigma)\right]^{m}.$ (12)
@@
$\lim_{m\to\infty}\frac{1}{m}\ln\left(\alpha(M_{m})+\beta_{\rho,\sigma}(M_{m})\right)=-\tilde{D}_{1/2}(\rho\|\sigma),$ (13)
where $\tilde{D}_{1/2}(\rho\|\sigma):=-\ln(F(\rho,\sigma))$ is the sandwiched Rényi Relative Entropy [37], [38] between $\rho$ and $\sigma$ of order $1/2$.
In Section A-A we prove Theorem 1 and in Section A-B we replicate a result from [36] which demonstrates our result is tight in the sense that for any other decision rule $\tilde{M}_{m}$ we will have
$\liminf\limits_{m\to\infty}\frac{1}{m}\ln\left(\alpha(\tilde{M}_{m})+\beta_{\rho,\sigma}(\tilde{M}_{m})\right)\geq-\tilde{D}_{1/2}(\rho\|\sigma).$ (14)
@@ III-B Qubit QUHT
Next, we provide an example illustrating how leveraging a concentration result for a quantum state estimator allows us to derive a corresponding result in the QUHT framework. Let $I_{2}$ be the $2\times 2$ identity and $\sigma_{X}=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$, $\sigma_{Y}=\begin{pmatrix}0&-i\\ i&0\end{pmatrix}$, $\sigma_{Z}=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}$ the Pauli matrices, and matrix $\vec{\sigma}=[\sigma_{X},\sigma_{Y},\sigma_{Z}]^{T}$. Any Hermitian $A\in L(\mathcal{H}_{2})$ admits the Bloch decomposition
$A=\frac{1}{2}\bigl{(}I_{2}+\vec{r}\!\cdot\!\vec{\sigma}\bigr{)}$ (15)
Where $\vec{r}=[\operatorname{Tr}[A\sigma_{X}],\operatorname{Tr}[A\sigma_{Y}],\operatorname{Tr}[A\sigma_{Z}]]$
[…]
Let $\sigma\in\mathcal{D}(\mathcal{H}_{2})$ and $m$ be divisible by three. Denote the joint state of $m$ copies by $\sigma^{\otimes m}$, and partition its factors into three sets $G_{i}$ ($i\in\{X,Y,Z\}$) of size $m/3$. On each $G_{i}$, perform the $0,1$-valued projective measurement $M_{i}$ in the eigen-basis of the Pauli operator $\sigma_{i}\in\{\sigma_{X},\sigma_{Y},\sigma_{Z}\}$. The empirical Bloch coordinate is $\hat{r}_{i}=\frac{3}{m}\sum_{\ell\in G_{i}}M_{i}(\sigma)$, which satisfies $\mathbb{E}[\hat{r}_{i}]=\operatorname{Tr}[\sigma_{i}\sigma]$. The reconstructed state is
$\hat{\sigma}=\tfrac{1}{2}\Bigl{(}I+\sum_{i\in\{x,y,z\}}\sigma_{i}\,\hat{r}_{i}\Bigr{)}.$ (16)
@@
$\mathbb{P}\left[\left\|\sigma-\hat{\sigma}\right\|_{1}>\varepsilon\right]\leq 6\exp\left\{-\frac{m\varepsilon^{2}}{54}\right\}.$ (17)
As the number of samples increases, the confidence region around $\sigma$ becomes more concentrated, with $\rho$ eventually residing outside it, as seen in Figure 1. The proof of this result can be found in Section A-C.
@@
$\beta_{\rho,\sigma}(M_{m})$ $\lesssim\exp\left\{\frac{-m\left\|\rho-\sigma\right\|_{1}^{2}}{54}\right\},$ (19)
where the $\lesssim$ hides an additive $O\left(\left\|\rho-\sigma\right\|_{1}m^{1/2}\log^{1/2}(1/\alpha)\right)$ term in the exponent.
The proof can be found in Section A-D.
We can even use this concentration result to define a strategy for (mixed) two-sample qubit QUHT.
@@
$\beta_{\rho,\sigma}(M_{k})$ $\lesssim\exp\left\{-\frac{k\left\|\rho-\sigma\right\|_{1}^{2}}{216}\right\},$ (21)
where $\lesssim$ is hides an additive $O(k^{1/2}\ln(1/\alpha))$ term in the exponent.
The proof can be found in Section A-E.

### v3

@@ I Introduction
This parallel - both in problem formulations and in solutions - between classical and QHT is quite apparent in many of the results in the literature. Indeed, this observation is what motivates the present work: although there is a line of work focused on it in classical hypothesis testing, the so-called universal hypothesis testing (UHT) has remained largely unexplored in the quantum setting. UHT addresses the problem of determining whether a sequence of samples is generated from a known nominal distribution. It is, in essence, equivalent to the goodness of fit testing [7, 6, 8, 9]. Hoeffding’s formulation for UHT [10], in which the unknown alternative is a fixed distribution in the probability simplex, leads to a surprising result: even though this alternative is completely unknown, the achievable performance (in terms of the type II error probability, or more precisely, its exponential decay rate) is essentially the same as that of simple hypothesis testing when the alternative is fully specified.
@@
While QUHT has not been formally defined or studied to date, there have been several related works reported in the quantum literature. These include extensions of Sanov’s theorem [11, 12, 13, 14], exentions of classical $t$ and $F$ tests for Gaussian quantum states [15], composite hypothesis testing [16, 17, 18], and Hoeffding-type error exponents in binary quantum testing [19, 20, 12]. Additionally, there are several lines of research in quantum circuits and information where QUHT can play an important role. Examples include quantum circuit equivalence [21, 22], quantum state discrimination [23, 24, 25, 26], and quantum state verification and certification [27, 28, 29].
The UHT formulation described above is often referred to as the one-sample test, especially in the machine learning (ML) community. One can similarly define the quantum analog to the two-sample test [30] in which two sequences of samples (quantum systems) are given and the task is to determine whether the two sequences are generated from the same distribution (quantum state). Surprisingly, even in the classical setting the optimal asymptotic error exponent of two sample testing was only recently shown, specifically in the continuous distribution setting [31] [32]. Our proposed solution to the QUHT can be easily generalized to the two-sample setting, since respective concentrations of the two-sample sequences will ensure exponential consistency of the hypothesis test when QST is applied to both sequences.
@@ II Problem Setup
Let $\mathcal{H}$ be a finite‑dimensional Hilbert space. We denote by $\mathcal{L}(\mathcal{H})$ the space of linear operators on $\mathcal{H}$ and by $\mathcal{U}(\mathcal{H})\subset\mathcal{L}(\mathcal{H})$ the unitary operators. The set of density operators $\mathcal{D}(\mathcal{H})\subset\mathcal{L}(\mathcal{H})$ consists of positive semidefinite, unit‑trace operators [33]. In the special case $\dim(\mathcal{H}_{2})=2$, $\mathcal{D}(\mathcal{H}_{2})$ is the set of qubit density operators. We write $\operatorname{Tr}(\cdot)$ for the trace functional, and for any $\sigma\in\mathcal{D}(\mathcal{H})$, its $m$-fold tensor product is denoted $\sigma^{\otimes m}\in\mathcal{D}(\mathcal{H}^{\otimes m})$. Throughout, $\ln$ denotes the logarithm base $e$. The trace norm of $\rho\in\mathcal{L}(\mathcal{H})$ is $\|\rho\|_{1}=\operatorname{Tr}[\sqrt{\rho^{\dagger}\rho}]$, and the trace distance between $\rho$ and $\sigma$ is $T(\rho,\sigma)=\tfrac{1}{2}\|\rho-\sigma\|_{1}$. The quantum fidelity is $F(\rho,\sigma)=\|\sqrt{\rho}\,\sqrt{\sigma}\|_{1}^{2}$, and the quantum relative entropy is $D(\rho\|\sigma)=\operatorname{Tr}[\rho\ln\rho-\rho\ln\sigma]$ [33]. We use $D_{KL}(P\|Q)$ to denote the classical Kullback-Leibler divergence.
@@ II-A Quantum Hypothesis Testing
$\begin{split}H_{0}:\sigma&=\rho_{0}\\ H_{1}:\sigma&=\rho_{1}.\end{split}$ (1)
Any test is specified by a decision rule $M_{m}:\mathcal{D}(\mathcal{H})^{m}\to\{0,1\}$, such that
$\hat{H}=\begin{cases}H_{0},&\text{if }\ M_{m}(\sigma^{\otimes m})=0\\ H_{1},&\text{if }\ M_{m}(\sigma^{\otimes m})=1\end{cases}$ (2)
Here we use $\hat{H}$ to denote the declared hypothesis and will reserve $H$ for the true hypothesis.
For the decision rule described by $M_{m}$ define the type I error to be $\alpha(M_{m}):=\mathbb{P}_{M_{m}}[\hat{H}=H_{1}|H=H_{0}]$ and the type II error to be $\beta(M_{m}):=\mathbb{P}_{M_{m}}[\hat{H}=H_{0}|H=H_{1}]$. The goal of any good testing strategy would be to minimize these errors, with respect to some constraints. Two common paradigms for developing tests are the so called “symmetric” and “asymmetric” [34] hypothesis testing settings. In the symmetric setting the goal is to simply minimize the total error or, when prior probabilities are available, the average error probability, whereas in the asymmetric setting, the goal is to minimize the type II error subject to a constraint on the Type I error. In this work we will be primarily concerned with the asymmetric setting, with minor attention paid to the symmetric setting. In classical literature the symmetric setting is sometimes referred to as the Bayesian setting, and the asymmetric setting is classically known as the Neyman-Pearson setting.
With a single (or finite) copy of the quantum systems, the Helstrom bound [35] [2] provides a lower bound on the symmetric error probability (which is in essence the average error probability, if an equal prior is assumed):
$\alpha(M_{m})+\beta(M_{m})\geq\frac{1}{2}\left(1-\frac{1}{2}\left\|\rho_{0}^{\otimes m}-\rho_{1}^{\otimes m}\right\|_{1}\right).$ (3)
The Quantum Stein’s lemma [4, 5], on the other hand, provides an asymptotic performance bound for the asymmetric setting when a large number of copies are available. Specifically the lemma states that for all $\varepsilon\in(0,1)$,
$\lim_{m\rightarrow\infty}\frac{1}{m}\log\beta^{*}(\epsilon)=-D(\rho_{0}\|\rho_{1}),$ (4)
where
$\beta^{*}(\epsilon)=\min_{M_{m}:\alpha(M_{m})<\epsilon}\beta(M_{m}).$ (5)
@@ II-B Quantum Universal Hypothesis Testing
Before describing the QUHT problem, we find it helpful to revisit its classical equivalent. Given samples $X^{m}\overset{\text{i.i.d.}}{\sim}Q$ from an unknown distribution, the goal is to determine if the true distribution is a known nominal distribution $P$ or not, i.e.
$\begin{split}H_{0}:Q=P\\ H_{1}:Q\neq P.\end{split}$ (6)
This problem was considered in a seminal work by Hoeffding [10], where he showed that the classical relative entropy, $D_{KL}(P\|Q)$, was, surprisingly, also the optimal achievable type II error exponent, when the type I error decayed in a bounded manner. One subtle note about his work was his consideration of the error under the alternative hypothesis, as he treated the true distribution as unknown but fixed.
QUHT is a natural extension of the classical UHT framework to the quantum setting. Given an $m$-fold tensor product state $\sigma^{\otimes m}$, the goal is to determine if $\sigma$ is a known nominal state, $\rho$, i.e.
$\begin{split}H_{0}:\sigma&=\rho\\ H_{1}:\sigma&\neq\rho.\end{split}$ (7)
The QUHT formulation directly mirrors Hoeffding’s formulation of the classical UHT, where the alternative is represented by an unknown but fixed state $\sigma$ that is different from $\rho$. For a given decision rule $M_{m}$, we denote the type I and II error probabilities as $\alpha(M_{m})$ and $\beta(M_{m})$, respectively. We note that while any decision rule should not depend on $\sigma$, as it is unknown apriori, the actual performance (in particular, type II error probability) is typically a function of $\sigma$. See ˜1a for a visual representation.
The merit of the QUHT formulation lies in the generality of the alternative hypothesis. While this may motivate its use in broad applications of quantum systems, it also presents unique challenges. Many techniques developed for QHT rely on knowledge of both states — e.g., optimal tests for structured alternatives [36, 11, 13] - and are thus not directly applicable when the alternative is entirely unspecified. An exception is found in recent work [14], which constructs a "quantum empirical distribution" and proves large-deviation results reminiscent of those in the classical setting. We further discuss its connection to our work in Section IV.
@@ II-C Quantum Two-Sample Test
$\begin{split}H_{0}:\sigma&=\rho\\ H_{1}:\sigma&\neq\rho.\end{split}$ (8)
While the form appears identical to the one-sample test (i.e., QUHT), the setup differs in one key way. The two-sample test has no knowledge of both $\rho$ and $\sigma$ and is given two sequences of samples, whereas the one-sample test has perfect knowledge of $\rho$ and uses a single sequence generated from the unknown state $\sigma$. See ˜1b for a visual of the problem and our proposed solution.
Figure 1a QUHT One Sample Problem. The yellow circle denotes the confidence region around our estimate $\hat{\sigma}$, with this region shrinking exponentially as $n$ increases. When $\rho$ is within this confidence region, we accept $H_{0}$.
Figure 1b QUHT Two Sample Problem. Colored circles represent the estimated states of $\rho$ and $\sigma$, while the enclosing balls denote the confidence regions surrounding the estimates. Depending on the overlap between the confidence regions we may accept or reject $H_{0}$.
@@ III Results
This section presents our main results on QUHT for both the one- and two-sample settings. As alluded to in Section I, we primarily leverage the concentration results found in the QST literature to obtain the error exponent for type II error probabilities. The only exception is when the nominal state is pure, in which case we obtain an optimal error exponent. Proofs of all results can be found in the appendices of this paper.
Table˜I summarizes our main results for both one‑sample and two‑sample variants of QUHT. To emphasize the dominant error‑exponent terms, we have omitted those that vanish asymptotically as the number of copies of the quantum states under test grows without bound. Here, $m$ and $n$ denote the numbers of identical copies of the states $\rho$ and $\sigma$, respectively, as introduced in Section II-C. Throughout our analysis, the type I error threshold $\alpha\in(0,1)$ is treated as a fixed constant. In the two‑sample setting, we restrict attention to the regime in which $m$ and $n$ increase at the same rate and define $k=\min\{m,n\}$ to express all bounds in terms of the effective sample size.
@@
Theorem˜1
Indep. $F(\rho,\sigma)]^{m}$
@@
Theorem˜2
Indep. Lemma˜1 $\exp\left\{-\frac{m\left\|\rho-\sigma\right\|_{1}^{2}}{54}\right\}$
@@
Theorem˜3
Indep. Lemma˜1 $\exp\left\{-\frac{k\left\|\rho-\sigma\right\|_{1}^{2}}{216}\right\}$
@@
Theorem˜4
Indep. Lemma˜2 $\exp\left\{-\frac{m\left\|\rho-\sigma\right\|_{1}^{2}}{86\cdot d^{3}}\right\}$
@@
Theorem˜5
Indep. Lemma˜2 $\exp\left\{-\frac{k\left\|\rho-\sigma\right\|_{1}^{2}}{344\cdot d^{3}}\right\}$
@@
Theorem˜6
Entangled Lemma˜3 $\exp\left\{-\frac{m\left\|\rho-\sigma\right\|_{1}^{2}}{2}\right\}$
@@
Theorem˜7
Entangled Lemma˜3 $\exp\left\{-\frac{k\left\|\rho-\sigma\right\|_{1}^{2}}{8}\right\}$
@@ III-A Pure State QUHT
Take $\rho=\lvert\phi\rangle\langle\phi\rvert\in\mathcal{D}(\mathcal{H})$ a pure state density operator and $\sigma\in\mathcal{D}(\mathcal{H})$ a (potentially mixed) density operator. In the QUHT setting, given $\sigma^{\otimes m}$, there exists a decision rule $M_{m}$ with independent measurements, such that
@@
$\beta(M_{m})$ $=\left[F(\rho,\sigma)\right]^{m}.$ (10)
@@
$\lim_{m\to\infty}\frac{1}{m}\ln\left(\alpha(M_{m})+\beta(M_{m})\right)=-\ln F(\rho,\sigma).$ (11)
In Section˜A-C we prove Theorem˜1 and in Section˜A-D we replicate a result from [34] which demonstrates our method is optimal in the sense that for any other decision rule $\tilde{M}_{m}$ we will have
$\liminf\limits_{m\to\infty}\frac{1}{m}\ln\left(\alpha(\tilde{M}_{m})+\beta(\tilde{M}_{m})\right)\geq-\ln F(\rho,\sigma).$ (12)
Note that $\tilde{D}_{1/2}(\rho\|\sigma):=-\ln(F(\rho,\sigma))$ is the sandwiched Rényi Relative Entropy [38], [39] between $\rho$ and $\sigma$ of order $1/2$.
@@ III-B Qubit QUHT
Next, we provide an example illustrating how leveraging a concentration result for a quantum state estimator allows us to derive a corresponding result in the QUHT framework. Let $I_{2}$ be the two dimensional identity matrix, and $\sigma_{X},\sigma_{Y},\sigma_{Z}$ be the standard Pauli matrices [40], and matrix $\vec{\sigma}=[\sigma_{X},\sigma_{Y},\sigma_{Z}]^{T}$. Any Hermitian $A\in L(\mathcal{H}_{2})$ admits the Bloch decomposition
$A=\frac{1}{2}\bigl(I_{2}+\vec{r}\!\cdot\!\vec{\sigma}\bigr),$ (13)
where $\vec{r}=[\operatorname{Tr}[A\sigma_{X}],\operatorname{Tr}[A\sigma_{Y}],\operatorname{Tr}[A\sigma_{Z}]]$.
Let $\sigma\in\mathcal{D}(\mathcal{H}_{2})$ and $m$ be divisible by three. Denote the joint state of $m$ copies by $\sigma^{\otimes m}$, and partition its factors into three sets $G_{i}$ ($i\in\{X,Y,Z\}$) of size $m/3$. On each $G_{i}$, perform the $0,1$-valued projective measurement $M_{i}$ in the eigen-basis of the Pauli operator $\sigma_{i}\in\{\sigma_{X},\sigma_{Y},\sigma_{Z}\}$. The empirical Bloch coordinate is $\hat{r}_{i}=\frac{3}{m}\sum_{\ell\in G_{i}}M_{i}(\sigma)$, which satisfies $\mathbb{E}[\hat{r}_{i}]=\operatorname{Tr}[\sigma_{i}\sigma]$, and define the vector $\hat{\vec{r}}=[\hat{r}_{x},\hat{r}_{y},\hat{r}_{z}]$. The reconstructed state is
$\hat{\sigma}=\tfrac{1}{2}\Bigl(I_{2}+\hat{\vec{r}}\!\cdot\!\vec{\sigma}\Bigr).$ (14)
@@
$\mathbb{P}\left[\left\|\sigma-\hat{\sigma}\right\|_{1}>\varepsilon\right]\leq 6\exp\left\{-\frac{m\varepsilon^{2}}{54}\right\}$ (15)
As the number of samples increases, the confidence region around $\sigma$ becomes more concentrated, with $\rho$ eventually residing outside it.
The proof of this result can be found in Section˜A-E.
@@
$\beta(M_{m})$ $\lesssim\exp\left\{\frac{-m\left\|\rho-\sigma\right\|_{1}^{2}}{54}\right\},$ (17)
where the $\lesssim$ hides an additive $O\left(m^{1/2}\right)$ term in the exponent.
We can use this concentration result to define a strategy for (mixed) two-sample qubit QUHT.
@@
$\beta(M_{k})$ $\lesssim\exp\left\{-\frac{k\left\|\rho-\sigma\right\|_{1}^{2}}{216}\right\},$ (19)
where $\lesssim$ hides an additive $O(k^{1/2})$ term in the exponent.